I need help solving this problem:
3^√x^2 = 4
"^" it means it's an exponent.... so it is 3 cubed to the square root of X squared = 4"

Answer by stanbon(75887)

I need help solving this problem:
3^√x^2 = 4
"^" it means it's an exponent.... so it is 3 cubed to the square root of X squared = 4"
------------
It looks like you have posted two problems:
1st: 3^(sqrt(x^2) = 4
3^x = 4
Take the log to get:
x(log3) = log4
x= (log4)/(log3)
x= 1.262
----------------
2nd: 3 cubed to the square root of X squared = 4"
(3^3)^(sqrt(x^2)) = 4
27^x = 4
Take the log of both sides to get:
xlog27=log4
x=(log4)/(log27)
x=0.42062
------------------
Cheers,
Stan